# Photovoltaic power forecasting using quantum machine learning

### v1

@@ II.5 Training and results
where $N$ is number of predictions, $x$ denotes the predicted PV power, and $y$ represents the actual PV power value.
@@
Here $\vec{x}=\left(x_{1},x_{2},\dots,x_{N}\right)$ and $\vec{y}=\left(y_{1},y_{2},\dots,y_{N}\right)$ represent vectors of predicted and target PV power values respectively, where $N$ is the number of predicted values.
@@
Table 1: The table shows which hyperparameters are being optimized, the limits of change, and the best values found during hyperparameter optimization.
@@
Table 2: Summary of the results for the proposed models. In a direct comparison, the HQNN outperforms the MLP in both training and testing losses, as evidenced across three critical metrics: RMSE, MAE, and MSE. Notably, the HQNN’s MSE is $41\%$ less than MLP’s one, which means that HQNN has more reliable forecasts, despite the fact that it has $1.8$ times fewer parameters. At the same time, the quality of HQLSTM’s on the test dataset is $40\%$ better compared to LTSM’s one on MAE and MSE metrics, and is $21\%$ better on RMSE metric, although the first one has less than half weights. In a broader comparison encompassing all four models, HQLSTM emerges as the most precise model on all metrics, namely on $52\%$ more precise than HQNN having two times fewer trainable parameters.
@@ II.5.1 HQNN & MLP and HQLSTM & LSTM
Both the HQNN and its classical analog, the MLP, were trained for $20$ epochs. In contrast, the HQLSTM and its classical counterpart, the LSTM, were trained for over $50$ epochs. The Adam optimiser kingma2017adam from the PyTorch framework was used to update the parameters of the models in order to minimize their loss functions. The comprehensive training process, accompanied by the results, is delineated in Fig. 4 and in Table 2.
@@
Furthermore, we opted to partition the dataset into training and testing sets in a $4:1$ ratio. This strategy was implemented to promote a comprehensive evaluation of our models, as we carried out model training and assessment on five distinct data splits. Subsequently, a meticulous averaging process was employed to consolidate the results obtained from these splits, and the mean values thus derived served as the primary metrics for inter-model comparisons.
The utilization of cross-validation techniques in our methodology significantly bolsters the robustness and reliability of our results, as they diminish the reliance on specific train-test partitioning, thereby enhancing the credibility of our findings.
In a head-to-head comparison between HQNN and MLP, the former exhibits superior performance regarding training and testing losses across three key metrics: RMSE, MAE, and MSE. Specifically, HQNN surpasses MLP’s power prediction accuracy by $41\%$ estimated by MSE loss function and by $26\%$ by MAE loss, all the while boasting $1.8$ times fewer parameters (2266& 3987226639872266\ \&\ 39872266 & 3987).
On juxtaposing HQLSTM with LSTM, the former outperforms in training and testing loss across all three aforementioned metrics. Remarkably, HQLSTM’s has better predictive ability (on $40\%$ better) than LSTM assessed by the MSE and MAE metrics, and it achieves this with less than half the number of parameters (1109& 2857110928571109\ \&\ 28571109 & 2857). Moreover, HQLSTMs are more resistant to overfitting, while classical LSTM suffers from it.
In a broader comparison encompassing all four models, HQLSTM emerges as the most precise model on all metrics, namely on $52\%$ more precise than HQNN, having two times fewer trainable parameters.
It is worth noting that we performed hyperparameter optimization technique using the Optuna optimizer optuna_link . The set of optimized parameters, limits of their variation, and best sets of hyperparameters for all our $4$ models are presented in Table 1.
Moreover, we scanned external articles that refer to this dataset and found only one article that solves $1$ hour ahead PV power prediction using neural networks. In comparison, our HQNN model is better than the model from the external article according to VAF metrics by $40\%$ (91& 65916591\,\&\,6591 & 65).
Further, to confirm that hybrid models train better including on a smaller dataset, an additional experiment was conducted, wherein the volume of training data was intentionally reduced. The results are shown in Figure 4. The hybrid models performed better with less data, having fewer losses and better prediction capabilities than classical models.
@@ II.5.2 HQSeq2Seq & Seq2Seq
After preprocessing the dataset, which is described in the Section II.1, it spanned $12775$ hours from 3/5/12 4:55 AM to 8/19/13 10:00 AM for training, and $3194$ hours from 8/19/13 11:00 PM to 12/30/13 00:00 AM for testing.
Although models are capable of being trained on sequences of arbitrary lengths, we chose to use sequences of fixed $96$ hours for simplicity. In this case, the encoder gets $96$ hours of all available features, while the decoder is asked to extrapolate only the “Power” feature of the data $96$ hours ahead. Training for $15$ epochs with the Adam optimizer (learning rate $0.001$) proved to be enough for the models to converge (Fig. 4 (f-g)). As an example of inference, we pass the time series of the length different from $96$ into both models and prompt them to give us a forecast for $137$ hours ahead (Fig. 4 (h)). We can conclude that both models transition from the fixed sequence length to an arbitrary one quite well. It may even be possible to improve these results by introducing variable-length sequences into the training stage. We also measured the dependency of test loss on the size of the training dataset for Seq2Seq and HQSeq2Seq, shown in Fig. 4 (e). As one can see, the test RMSE loss of the hybrid model is less for any size of training data, which proves that the hybrid model shows an advantage over the classical model, including on a trimmed dataset.
@@ III Discussion
In this work we introduced three hybrid quantum approaches to time series prediction task. The first two models allow one to predict the power of solar panels for $1$ hour ahead, using weather features for the previous $24$ hours. The third model allows to predict a longer-term user-defined forecast, showcasing the versatility of our models for various planning tasks.
@@
The first approach is the HQNN, a combination of classical fully connected layers and quantum layer, which is a VVRQ, analog of classical fully-connected layer. We compared this hybrid model with its classical counterpart, MLP, and demonstrated that, even though HQNN has $1.8$ times less variational parameters, it has $41\%$ better predictive ability, estimated by MSE error.
@@
The second approach is HQLSTM, a hybrid quantum analogue of classical LSTM. Here, QDI layer is inserted into each gate of the LSTM cell. This approach provides a $40\%$ improvement in prediction using the MAE and MSE metrics comraped to its classical counterpart. Our proposed architecture is a unique combination of classical and quantum layers, which we believe to be a breakthrough in solving time series prediction tasks. Comparing HQNN and HQLSTM models, the second one was better on $52\%$ than HQNN having two times fewer weights.
@@
The third approach is hybrid Seq2Seq model, a classical Seq2Seq model, consisting of $2$ LSTMs with quantum layer at the end. This approach allows one to predict the PV power not only for an hour ahead, but for any number of hours ahead, without knowing the weather features in advance. The addition of the proposed QDI layer improves the accuracy of the predictions, reducing the MAE error by $16\%$ compared to a purely classical Seq2Seq model. Our proposed architecture is a unique combination of classical and quantum layers, which we believe to be a breakthrough in solving time series prediction tasks.
@@
Also, for all models, we conducted an additional experiment in which our models were trained on a reduced dataset, and confirmed that hybrid models have better learning capabilities and have less loss trained on any amount of dataset compared to their classical counterparts. This confirmation can serve as an excellent motivation to use hybrid networks for applications where data collection is a complex task.
@@
It is worth noting that all parameters are trainable in our layers; the architecture and hyperparameters were selected by the Optuna optimizer. Moreover, we compare our models to a paper that solved the same problem using the same dataset, and demonstrate that our best HQLSTM is $40\%$ more accurate in predicting power using VAF metric.
@@
To fully unlock the potential of HQNNs in time series prediction problem, further research and testing of models on other datasets is necessary. Also, the development of more efficient optimization VQC training and implementation methods, larger-scale quantum hardware could lead to even more significant performance improvements.
@@
Furthermore, while this work was done on a public dataset with an emphasis on hybrid quantum models for better forecasting performance, the quality and source of data plays a crucial role in overall effectiveness in the real world, especially considering weather data. Accurate weather forecast is a crucial input into any high performing and useful PV prediction given its dynamism and influence on PV output. An interesting area of research is cloud prediction using satellite and weather data for geo-locations, directly impacting solar irradiance and therefore PV output. The added complexity could enhance the need for hybrid quantum models to increase computational efficiency and higher quality forecasts.
@@
To summarize, our developments provide three hybrid quantum approaches for time series problems that demonstrate the possibility of combining classical and quantum methods. Our proposed models show improved performance compared to classical models with similar architecture when using fewer variation parameters. We believe that these results pave the way for further research in developing hybrid models that leverage the strengths of both classical and quantum computing.

### v2

@@ II.5 Training and results
where $N$ is number of predictions, $\vec{x}=\left(x_{1},x_{2},\dots,x_{N}\right)$ is the predicted PV power, and $\vec{y}=\left(y_{1},y_{2},\dots,y_{N}\right)$ is the actual PV power value.
@@
Table 1: The table summarizes the hyperparameter optimization process. It shows which hyperparameters are being optimized, the range over which they could be varied, and the best values found.
@@
Table 2: Summary of the results for the proposed models. In a direct comparison, the HQNN outperforms the MLP in both training and testing losses, as evidenced across three critical metrics: RMSE, MAE, and MSE. Notably, the HQNN’s MSE is $41\%$ lower than that of the MLP. This demonstrates that the HQNN has more reliable forecasts, despite the fact that it has $1.8$ times fewer parameters. On the test dataset, the HQLSTM is $40\%$ better compared to the LTSM on the MAE and MSE metrics. It is also $21\%$ better on the RMSE metric, even though it has less than half as many weights. Overall, the HQLSTM emerges as the most precise model on all metrics. Specifically, it is $52\%$ more precise than the HQNN while having two times as few trainable parameters.
@@ II.5.1 HQNN & MLP and HQLSTM & LSTM
Both the HQNN and its classical analog, the MLP, were trained for $20$ epochs. In contrast, the HQLSTM and its classical counterpart, the LSTM, were trained for over $50$ epochs. The Adam optimiser kingma2017adam from the PyTorch framework was used to update the parameters of the models in order to minimize their loss functions. The comprehensive training process, accompanied by the results, is shown in Fig. 4 and in Table 2.
@@
Furthermore, we performed cross-validation by partitioning the dataset into training and testing sets in a $4:1$ ratio. This strategy was implemented to promote a comprehensive evaluation of our models, as we carried out model training and assessment on five distinct data splits. Subsequent averaging was used to consolidate the results obtained from these splits, and the mean values served as the primary metric for inter-model comparisons.
In a head-to-head comparison between the HQNN and the MLP, the former exhibits superior performance regarding training and testing losses across three key metrics: RMSE, MAE, and MSE. Specifically, the HQNN surpasses the MLP’s power prediction accuracy by $41\%$ as estimated by the MSE loss function, and by $26\%$ on the MAE loss, all while boasting $1.8$ times fewer parameters (2266& 3987226639872266\ \&\ 39872266 & 3987).
When comparing the HQLSTM and LSTM models, the former outperforms the latter in both training and testing losses across all three of the aforementioned metrics. Remarkably, the HQLSTM has a $40\%$ better predictive ability than the LSTM as assessed by the MSE and MAE metrics. It achieves this with less than half the number of parameters (1109& 2857110928571109\ \&\ 28571109 & 2857). Moreover, HQLSTMs are more resistant to overfitting, while classical LSTMs suffer from it.
In a broader comparison encompassing all four models, the HQLSTM emerges as the most precise model on all the metrics. Notably, it is $52\%$ more precise than the HQNN, while having half as many trainable parameters.
To achieve these results, we performed hyperparameter optimization using the Optuna optimizer optuna_link . The set of optimized parameters, the limits on their variation, and the best sets of hyperparameters for all our four models are presented in Table 1.
Moreover, we searched for external articles that refer to this dataset and found only one article that solves the $1$ hour ahead PV power prediction using neural networks kaloop2021novel . In comparison, our HQNN model is better than the model from this external article according to the VAF metric by $40\%$ (91& 65916591\,\&\,6591 & 65).
Further, to confirm that hybrid models train better including on a smaller dataset, an additional experiment was conducted where the volume of training data was intentionally reduced. The results are shown in Figure 4. The hybrid models performed better with less data, having lower losses and better prediction capabilities than the classical models.
@@ II.5.2 HQSeq2Seq & Seq2Seq
After the preprocessing steps described in the Section II.1, the dataset spanned $12775$ hours from 3/5/12 4:55 AM to 8/19/13 10:00 AM for training, and $3194$ hours from 8/19/13 11:00 PM to 12/30/13 00:00 AM for testing.
Although the models are capable of being trained on sequences of arbitrary lengths, we chose to use sequences with a fixed $96$ hour length for simplicity. In this case, the encoder gets $96$ hours of all available features, while the decoder is asked to extrapolate only the “Power” feature of the data $96$ hours ahead. Training for $15$ epochs with the Adam optimizer (learning rate $0.001$) proved to be enough for the models to converge (Fig. 4 (f-g)).
As an example of inference, we pass the time series of the length different from $96$ into both models and prompt them to give us a forecast for $137$ hours ahead (Fig. 4 (h)). We can conclude that both models transition from the fixed sequence length to an arbitrary one quite well. It may even be possible to improve these results by introducing variable-length sequences into the training stage.
We also investigated the dependence of the test loss on the size of the training dataset for Seq2Seq and HQSeq2Seq, as shown in Fig. 4(e). As one can see, the test RMSE loss of the hybrid model is less for any size of training data, showing that the hybrid model has an advantage over the classical model, including on a substantially trimmed dataset.
@@ III Discussion
In this work, we introduced three hybrid quantum approaches to time series prediction. The first two models allow one to predict the power of solar panels $1$ hour ahead using weather features from the previous $24$ hours. The third model is the most versatile since it enables longer-term, user-defined forecasts to be made.
@@
The first approach is the HQNN, a combination of classical fully connected layers and a VVRQ layer, which is a quantum analog of such layers. We compared this hybrid model with its classical counterpart, the MLP, and demonstrated that, even though the HQNN has $1.8$ times fewer variational parameters, its predictive ability is $41\%$ better, as estimated by the MSE metric.
@@
The second approach is the HQLSTM, a hybrid quantum analogue of the classical LSTM. Here, a QDI layer is inserted into each gate of the LSTM cell. This approach provides a $40\%$ improvement in prediction using the MAE and MSE metrics compared to its classical counterpart.
The HQLSTM was even an improvement on the HQNN despite having half as many weights.
@@
The third approach is the hybrid Seq2Seq model, which consists of two LSTMs with a quantum layer at the end. This approach allows one to predict the PV power not only for an hour ahead, but for any number of hours ahead, without knowing the weather features in advance. The addition of the proposed QDI layer improves the accuracy of the predictions, reducing the MAE error by $16\%$ compared to a purely classical Seq2Seq model.
@@
We conducted an additional experiment in which all our models were trained on a reduced dataset. It confirmed that hybrid models have better learning capabilities and lower losses than their classical counterparts when trained on any amount of data. This motivates the use of hybrid networks for applications where data collection is a complex task.
@@
We also compared our models to a paper that solved the same problem using the same dataset. We found that our best HQLSTM is $40\%$ more accurate in predicting power using the VAF metric.
@@
To fully unlock the potential of hybrid quantum approaches in time series prediction problems, further research and testing of models on other datasets is needed. More efficient optimization and training of VQCs, combined with higher fidelity and larger-scale quantum hardware, could lead to significantly greater performance improvements.
@@
While this work was done on a public dataset with an emphasis on hybrid quantum models for better forecasting performance, the quality and source of data plays a crucial role in overall effectiveness in the real world, especially considering weather data. Accurate weather forecast is a crucial input into any high performing and useful PV prediction given its dynamism and influence on PV output. An interesting area of research is cloud prediction using satellite and weather data for geolocations directly impacting solar irradiance and therefore PV output. The added complexity could enhance the need for hybrid quantum models to increase computational efficiency and higher quality forecasts.
@@
To summarize, our developments provide three hybrid quantum approaches for time series problems that demonstrate the possibility of combining classical and quantum methods. Our proposed models show improved performance compared to classical models with similar architectures when using fewer variational parameters. We believe that these results pave the way for further research in developing hybrid models that leverage the strengths of both classical and quantum computing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability All datasets used in this study are publicly available and are cited within the Dataset subsection of the paper. The machine learning models employed are fully described in the Results section, enabling independent reproduction of the results.